Sol Summary – March 16th

Crew 190 Sol Summary Report
16March2018
Sol 5

Summary Title:
Phoenix awaken! But still a bit buzzy.

Michael Saint-Guillain (XO)

Mission Status:
Ready to continue simulation on Sol 6, 08:00

Sol Activity Summary:
8:00 Gym
8:15 Medic inspection: memory and reflex tests
8:40 Breakfast
9:30 EVA: Bastien (cartography – 3D mapping), Martin (experiment check), Sophie (muography), Fred (photos)
9:30 Others: Scientific work
11:00 EVA ended; Crew meeting before lunch
12:30 Lunch
14:30 Scientific work: muography (Sophie), video extraction & compilation (Bastien), chemistry/biology (Ariane, Martin, Maximilien, Frederic), GreenHab (Mario), Phoenix troubleshooting (Michael)
18:00 Report redaction
19:00 CapCom

Look Ahead Plan:
Sol 6:

Notes: EVA (Max, Mario, Ariane, Mick) in the morning (scientific work for the others). Scientific work in the afternoon.

Anomalies in work:
Today we tried to get the Phoenix remote controlled rover working; after fixing some battery issues, we noticed that the rover was not able to connect to the (dedicated) router wifi. We had to manually change the clock of the rover it self in order to fix that problem: the rover had the correct date, but not the router, which we cannot fix (probably the internal clock battery of the router is dead). With similar date/times, router and rover are now able to talk to each other. The video transmission is working, but the commands are still bugging (can’t move the wheels, and cams movement is working randomly).

Weather:
Sunny in the morning. Cloudy in the afternoon. Nice weather.

Crew Physical Status:
Good.

EVA:
See EVA report.

Reports to be file:
Commander report
Journalist report
EVA#7 request for Sol6
Engineer report
GreenHab report
HSO report
Pictures

Support Requested:
None

Journalist Report – March 16th

The MDRS’ GreenHab is an extremely lush, green place. Filled with large amounts of tomato plants, as well as many other plants in smaller amounts, it is an oasis of life in the middle of the Martian desert. The tomatoes have been yielding generously – currently over a hundred cherry tomatoes are on the path of ripening, soon to be ready to be enjoyed by the crew, and other vegetables and leafy greens are growing well. However, all the plants are growing in soil, a medium that has had to be brought over from earth at large expenses. Our two botanists, Frederic and Mario, are both working on improving the disadvantages that come from soil as a medium for growth. On the one hand, Frederic is focusing on growing plants in Martian regolith (the dusty, iron-rich sands that make up most of Mars’ surface and give it its red colour). His research focuses on the advantages of different types of regolith, mixed in various amounts, with or without HydroGel. HydroGel is a man-made substance, an acrylate polymer, which has the incredible propriety of being able to absorb over 100 times its own weight in water, and release it slowly. This is extremely useful in dry climates back on Earth, as well as on Mars, where water is extremely precious.

On the other hand, Mario is working on a project that is somewhat the opposite : a hydroponics system. In such a system, plants are growing without soil, or any nutritive medium. Instead, their roots are hanging in a recirculated flow of nutrient-rich water, feeding them without wasting any water. In such a system, the nutrient composition of the water, as well as its pH and other factors, can be controlled with great precision to suit the plants. In addition to this, no water is wasted since it is recirculated, and the vertical format of the towers allows for minimal waste of space, also a precious resource in an extra-terrestrial station.

Both experiments are well underway! Fred has finished planting all his seeds and is waiting for them to sprout to start taking measures, whereas Mario has finished building the last of his three towers today.

All other experiments are also well underway ! Time is passing fast, and we are soon approaching the middle of our stay in the MDRS, even though we feel like we just got here. On Sunday, as a middle mark of our stay, we are planning a break day in our experiments, and a small celebration (with more cake!). In the meanwhile, we’ll keep on working hard, as there is no shortage of things to do here on Mars.

Operations Report – March 16th

Hello CapCom,

Here is my sixth engineering report, feel free to give me feedback:

Crew 190 Operations Report 16th March 2018

SOL: 5

Name of person filing report: Bastien BAIX

Non-nominal systems: –

Notes on non-nominal systems: –

Generator (hours run): Turned off at 8.30 am and turned on at 6.45 pm.

Solar – SOC 95% (Before generator is run at night)

Diesel – ~45%

Propane – 81%

Ethanol Free Gasoline (5 Gallon containers for ATV) – 2 empty jerrycans.

Water (trailer) – 0 gallons.

Water (static) – ~ 250 gallons

Water (third tank) – 500 gallons

Trailer to Static Pump used – NO

Water (loft) – Static to Loft Pump used – YES

Water Meter: 132 251.9

Toilet tank emptied: YES

ATV’s Used: –

Oil Added: NO

ATV Fuel Used: NO

Hours the ATVs were used today: _

Notes on ATVs: –

Deimos rover used: NO

Hours: 116.6

Beginning charge: 100%

Ending charge: –

Currently charging: NO

Notes on Deimos rover: flat tire, not yet repaired

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: NO

Hours: 24.8

Beginning charge: 100%

Ending charge: –

Currently charging: NO

Opportunity rover used: NO

Hours: 24.1

Beginning charge: 100%

Ending charge: –

Currently charging: YES

Curiosity rover used: NO

Hours: 18.1

Beginning charge: 100%

Ending charge: –

Currently charging: YES

HabCar used and why, where: –

General notes and comments: –

Summary of internet: ~ 200 Mb remaining

Summary of suits and radios: –

Summary of Hab operations: Ariane, Mario & Bastien working on their experiments

Summary of GreenHab operations: –

Summary of ScienceDome operations: Martin, Maximilien & Fred working on their experiments.

Summary of RAMM operations: Sophie adjusting her muon detector

Summary of health and safety issues: –

Questions, concerns and requests to Mission Support: –

Enjoy your evening,

Cheers,

Bastien Baix

Crew 190 Engineer – UCL to Mars 2018

EVA Report – March 17th

Hi CapCom,

Here is the EVA Report #6 for SOL5,

Thank you very much.

EVA Report:

EVA #6 Crew members: Martin ROUMAIN (EVA Leader), Bastien BAIX, Sophie WUYCKENS, Frédéric PEYRUSSON.

Location:

– Near the Hab: 518200E-4250750N

– Near Pooh’s Corner: 518900E-4251000N

Time:

Departure at 09:31 a.m.

Back at 11:01 a.m.

Duration:

1 hour 30 minutes

Narrative:

Everything went just fine during this sixth EVA. We went out of the airlock at 9:31 a.m. It was a sunny morning.

We started the EVA by digging a small hole for Sophie’s muons detector near the RAM (hole is not deeper than 20 cm and 30 cm wide).

We then checked Martin’s samples near the solar panel: they looked just fine. We were on our way to Pooh’s Corner when we saw the Mars Desert Research Station panel on the ground. Gusts must have made it fall. We took the opportunity to stop and replace it.

After a few minutes walking on the main road, we saw a hill and thought that the view from the top would be perfect for pictures. We took some pictures of the station in the sun.

We went back on the road and arrived to Pooh’s Corner, where we climbed (with some difficulties I must admit) another, higher, hill. We shot some videos and then went back home where we had a few tasks left.

When near the Hab, Bastien started to fly his drone for the 3D-mapping of the area, while Sophie got busy installing her detector right next to the RAM. Frédéric and Martin went behind the station to pick up the non-burnable trash, following Dr. Rupert’s demand. They found a lot of cans, metal waste and even a halogen lamp, probably from another crew. They put all the trash in a bag and put it in the main airlock. They then helped Sophie finish her installation, while Bastien was still on his mapping.

When Sophie and Bastien finished their work, all crew members went in the airlock for pressurization. That was a nice and productive EVA.

Purposes:

Check Martin’s samples near the Hab

3D-Mapping with Bastien’s drone

Find a location for Sophie’s muons detector

Additional purpose:

Take care of the non-burnable trash behind the station, following Dr. Rupert’s demand

Vehicles: None

9:28 a.m.: All crew members in the main airlock, ready for depressurization.

9:31 a.m.: Beginning of the EVA, started to dig a little hole for Sophie’s muons detector near the RAM.

9:48 a.m.: On the road to Pooh’s Corner, checked Martin’s samples near the solar panels, replaced the MDRS panel which fell because of the wind.

9:56 a.m.: Climbed to the top of a hill near the Hab to take pictures and shoot videos with Bastien’s drone.

10:12 a.m.: Climbed another hill near Pooh’s Corner, took pictures and shot videos.

10:22 a.m.: On the road back to the Hab.

10:32 a.m.: Took care of the non-burnable trash behind the Hab, helped Sophie install her muons detector while Bastien started 3D-mapping the area around the Hab.

10:34 a.m.: Put the non-burnable trash in the airlock.

10:52 a.m.: Sophie finished installing her detector near the RAM.

10:58 a.m.: Bastien finished 3D-mapping the area around the Hab, all crew members back in the main airlock with the non-burnable trash.

11:01 a.m.: All crew members safely back in the Hab.